NANUC - Canada's NMR Resource
../NMR SchedulesApply for TimeContactSite Map
HomeAbout NANUCFeaturesFacilitiesNMRResearchNMR ResourcesNewsDownloads
Features
Features
Tech Tips
Newsletter
Spectrum of the Month
Jobs
   
 
Quick Links
NMR Schedules
Apply for Time
Contact
Site Map
Tech Tips

Manostat Troubleshooting and Replacement, July 2003
by Deryck Webb


Figure 1 - Initial periodic lock change
In late June a substantial fluctuation was noticed on the acqmeter of the NANUC 800MHz system. The lock level was rising and falling over a wide range (Figure 1). There was no experiment operating at the time so it was deduced that their must be a radical variation in either the temperature of the sample or the temperature or pressure of the lower He bath. The VT was at a constant temperature and according to the datalogger so was the temperature of the lower He bath. It was determined that the pressure of the upper bath and it's subsequent affect on the lower bath, which was found to be the problem.

The Manostat pressure level was lowered to 970 mbar and the lock initially seemed to stabilize during the pressure drop. Then when the pressure reached a new equilibrium, the lock cycling appeared again. Data was collected for the Manostat at 15 minute intervals and then at 1 minute intervals and it was found that the cycles were occurring more often than first thought (see Figure 2).


Figure 2 - Manostat readings using 15 min. and 1 min. data saves

On July 8 Oxford was contacted and it was suggested checking all upper bath venting joints for leaks. All joints were checked and confirmed sealed, with one Tygon coupling at the rear of the Manostat being replaced as a precaution. Also, the 2 psi emergency release valve on the Manostat was tested at 930 mbar with no leaks detected. It was then suggested raising the pressure of the Manostat to 20 mbar over the atmospheric pressure. The atmospheric pressure of the room was 963 mbar (taken from room barometer) and the Manostat was set to 985 mbar. The pressure reached a value of 983 mbar and then began oscillating between 989 mbar and 980 mbar with a fluctuating He flow observed from the top bath. It was decided a new Manostat would be needed and in the meantime the current Manostat would be disconnected and a one-way, 2 psi valve would be placed on the He manifold of the magnet (see Figure 3). The flow through this valve was then measured through the existing top bath flow meter. The new Manostat arrived on July 14, 2003 (see Figure 4).


Figure 3 - 2 psi one-way valve on He manifold

Figure 4 - New Manostat, July 14, 2003

The new Manostat was installed during the computer rebooting and He fill on July 14, 2003. It was set to 975 mbar and initially went up to 977 mbar. NANUC standard #7 was then shimmed to a lock level of ~40% and left overnight.

The Manostat regulated a constant pressure overnight at 976 mbar with the setting remaining at 975mbar. The datalogger read 978 mbar. Everything appeared stable and Oxford suggested the 2 mbar difference from the Manostat to the datalogger was due to the Manostat being calibrated at the factory, at sea level. Regular experiments on the Spectrometer resumed in the afternoon of July 15, 2003.



Tech Tips Archive
Uninterruptible Power Supplies
The Nanuc 500MHz Cold Probe
Manostat Troubleshooting and Replacement, July 2003
Transporting Your NMR Samples to NANUC
Oxygen Sensors
Liquid Nitrogen Fills
 © 2002 NANUC - nanuc_nmr@nanuc.ca 101 NANUC, University of Alberta, Edmonton, AB, Canada  T6G 2E1